Solve each of the following systems by graphing.
(10) 2x – y = 4
2x –y = 6
Solve each of the following systems by addition. If a unique solution does not exist, state whether the system is inconsistent or dependent.
(12) 2x +3y = 1
5x + 3y = 16

Answer by tanimachatterjee(60)

10(2x)-y=4
2x-y=6
20x -y=4 ---------------(1)
-2x+y=-6 ----------------(2)
=18x =-2
x =-2/18
x=-1/9
substitute the value of x in equ 2
-2x+y=6
-2(-1/9)+y= -6
+2/9+y =-6
y=-6-2/9
y=-56/9
y=6.22




y=6.66